Determine a reasonable molecular formula for a compound with a molecular weight of 113. Show your work. b. Report the degrees of unsaturation for your formula.
Answer #1

Step 1

The structure with molecular formula C6H5Cl is a reasonable structure and have approx molecular weight of 113.

Reasonable structure means the set of structures describes the delocalization of electrons in a polyatomic ion or a molecule.

Benzene is the molecule which shows resonance that have molecular weight 78 g/mol. Thus from the given molecular weight it is conclude that the compound would be the substituted benzene. The remaining molecular weight (113-78 = 35) shows the substituent present at the benzene is chlorine atom (atomic weight 35.5 g).

Thus the structure of the compound is,

:C1 C6H5C1 Chlorobenzene

Step 2

The chlorine atom is the electron withdrawing group that is the group that draws the electron density toward it self.

Resonating structures for chlorobenzene are:

:ci: 2:01: :C1 :CI (1) (II) (III) (IV) (V) Resonating structures of chlorobenzene

Step 3: Degree of unsaturation

(b)

Degree of unsaturation tells about the total 'number of ring" and the number of 'pi bonds' present in the molecule.

There is formula to calculate the degree of unsaturation from the molecular formula.

Degree of unsaturation 20 +2 + N -H-X 2

Where, C represents number of carbon atom,

N represents number of nitrogen atom,

H represents number of hydrogen atom and

X represents number of halogen.

The molecular formula of the compound is C6H5Cl that is C=6, H=5 and X=Cl=1.

Thus degree of unsaturation in the compound is,

2C + 2 + N -H-X Degree of unsaturation = 2 2 x 6+2+0 -5-1 2 = 4

The degree of unsaturation in the compound is 4.

Also from the structure it is determined that the compound contains 1 ring ad 3 pi bonds that is the degree of unsaturation is 4.

Kindly upvote please,

Thank you.